# Black Hole Waterfall: a unitary phenomenological model for black hole evaporation with Page curve

### v1

@@ 2.1 The model of [1]
For physical BHs, a simple argument due to Zurek and Thorne [37] can be used to estimate ${n_{p0}}$ for a BH of initial mass $M_{0}$. Let the BH be composed of ${n_{p0}}$ particles of fixed mass $m_{0}$ of Compton wavelength $\frac{\hbar}{m_{0}c}$ such that $M_{0}={n_{p0}}m_{0}$. These collection of particles of mass $m_{0}$ cannot fit into a size smaller than the Schwarzschild radius $r_{s}=2GM/c^{2}$ of the BH without creating a BH themselves. Thus, setting $\frac{\hbar}{m_{0}c}\sim\frac{2GM_{0}}{c^{2}}$ yields $m_{0}=\frac{1}{2}M_{PL}(\frac{M_{PL}}{M_{0}})$. Setting $M_{0}={n_{p0}}\,m_{0}$ yields ${n_{p0}}=\frac{1}{2}(\frac{M_{PL}}{m_{0}})^{2}\sim\frac{1}{2}(\frac{10^{19}GeV}{m_{0}c^{2}})^{2}$, or $(\frac{m_{0}}{M_{PL}})^{2}=\frac{1}{2\,{n_{p0}}}$. As the BH evaporates, we take its energy to be $E_{BH}=E_{p}(\tau)=\bar{n}_{p}(\tau)m_{0}\,c^{2}=M(\tau)c^{2}$, with $\bar{n}_{p}(0)={n_{p0}}$. The BH thermodynamic (coarse grained) entropy is then given by $S_{\textrm{\tiny Bekenstein-Hawking}}(\tau)=\frac{1}{4}(4\pi r^{2}_{s})=4\pi(\frac{M(\tau)}{M_{PL}})^{2}=4\pi\bar{n}_{p}^{2}(\tau)(\frac{m_{0}}{M_{PL}})^{2}=2\pi\frac{\bar{n}_{p}^{2}(\tau)}{{n_{p0}}}\sim E^{2}_{BH}$ [38], while the fine-grained von Neumann entropy is given by $S(\rho_{BH})\leavevmode\nobreak\ =\leavevmode\nobreak\ S(\rho_{p})\leavevmode\nobreak\ \equiv\leavevmode\nobreak\ S(\rho_{i,s})$ (with the last equality following since the composite state is pure).
@@
Figure 3: (left) Evolution of the average number or particles in the pump/BH $\bar{n}_{p}(\tau)$ (black dashed line), and the signal/Hawking radiation $\bar{n}_{s}(\tau)$ (black solid) for the BH in an initial coherent state with $\bar{n}_{p}(0)$ = 35, and the signal/idlers in the vacuum state. $d^{eff}_{p}\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr{\scriptsize\textrm{def}}\cr\kern 2.0pt\cr=\cr\kern-2.0pt\cr}}}1+\Delta\bar{n}_{p}$ (gray dashed) and $d^{eff}_{s}\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr{\scriptsize\textrm{def}}\cr\kern 2.0pt\cr=\cr\kern-2.0pt\cr}}}1+\Delta\bar{n}_{s}$ (gray solid) are the effective dimensions of the the pump/BH and signal/Hawking radiation in terms of the variance in their particle number [1, 30]. Vertical red-dashed line indicates the time at which $d\bar{n}_{p}(\tau)/d\tau=0$ again for $\tau>0$, the limit of the validity of the model. (right) von Neumann entropies for: $S(\rho_{s})$ signal/Hawking radiation (dashed-black curve), $S(\rho_{thermal})$ effective thermal state (solid black curve) formed from $\rho_{s}$. The difference between the solid-black and dashed-black curve is the Page Infomation $I(\tau)\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr{\scriptsize\textrm{def}}\cr\kern 2.0pt\cr=\cr\kern-2.0pt\cr}}}S(\rho_{thermal})-S(\rho_{s})$.
In Fig.(3)(left) we plot the evolution of the average number or particles $\bar{n}_{p}(\tau)$ in the pump/BH (black dashed curve), and in the signal/Hawking radiation $\bar{n}_{s}(\tau)$ (black solid), for the BH in an initial coherent state with $\bar{n}_{p}(0)=35=\alpha^{2}$ [40], and the signal/idlers in the vacuum state, $|\psi(0)\rangle=|\alpha\rangle_{p}|0\rangle_{i}|0\rangle_{s}$, with $|\alpha\rangle_{p}=e^{-\alpha^{2}/2}\sum_{m=0}^{\infty}\frac{\alpha^{m}}{\sqrt{m!}}|m\rangle_{p}$ [4, 5]. For early times both $\bar{n}_{p}(\tau)$ and $\bar{n}_{s}(\tau)$ have nearly zero slope, and for these time with $\bar{n}_{p}(\tau)\gg\bar{n}_{s}(\tau)$ the Hawking radiation is essentially composed of squeezed signal/idler pairs, and hence is thermal. We define an effective thermal state from $\rho_{s}$ by defining $\rho_{thermal}=\sum_{m=0}^{\infty}\frac{\bar{n}_{s}^{m}(\tau)}{(\bar{n}_{s}(\tau)+1)^{m+1}}|m\rangle_{s}\langle m|$. The Page Information [30, 1] $I(\tau)\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr{\scriptsize\textrm{def}}\cr\kern 2.0pt\cr=\cr\kern-2.0pt\cr}}}S(\rho_{thermal})-S(\rho_{s})$ represents the information that leaks out of the BH due to the correlations (entanglement) built up between the pump/BH and signal/Hawking radiation as the system evolves. This is shown in Fig.(3)(right).
@@
We will see in the third model considered in Section 3, constructed in order to address model objective (3), this issue can also be rectified by assuming that the interior Hawking particles can also create their own interior/exterior idler/signal Hawking pairs, now with half of each interior particle’s energy, i.e. $\omega_{p}/4$. With each newly created idler acting as its own “pump”, this cascading (waterfall) process can be carried out ad infinitum so that the total energy carried away by the collection of all the signals/Hawking radiation is ${n_{p0}}\,\omega_{p}\,\sum_{d=1}^{D}(\frac{1}{2})^{d}=(1-(\frac{1}{2})^{D})\,{n_{p0}}\,\omega_{p}\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr{\scriptstyle{{D\to\infty}}}\cr\kern 2.0pt\cr{=\;\;}\cr\kern-2.0pt\cr}}}{n_{p0}}\,\omega_{p}$, the initial mass of the BH (this “remnant mitigation” proposal was first suggested by the author in [43]). What remains inside the horizon are ${n_{p0}}$ idler particles of arbitrarily small vanishing energy $(\frac{1}{2})^{D}\,{n_{p0}}\,\omega_{p}\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr{\scriptstyle{{D\to\infty}}}\cr\kern 2.0pt\cr{=\;\;}\cr\kern-2.0pt\cr}}}0$. This phenomena of zero energy remaining in the internal (“soft-hair”) horizon modes of the BH was explored in a 2018 paper by Hotta, Nambu and Yamaguchi entitled Soft-hair-enhanced entanglement beyond the Page Curves in Black Hole evaporation Qubit model [3]. In this work, an infinite number of supertranslation charges appear to be able to store the whole quantum information of the absorbed matter forming the BH, arising from a near-horizon symmetry providing a massive degeneracy of the Hamiltonian (the approximately zero energy “soft hair” of the BH horizon). The impetuous for the BH Waterfall Model presented here in this work was inspired by this work, with a desire to develop a qualitatively similar model, but from a squeezing Hamiltonian perspective. Before we explore this model in Section 3, which we descriptively denote as the Black Hole Waterfall, we first review the second One Shot model [2] in the next section, on which the former is built.
@@ 2.2 The One Shot model of [2]
The second model [2] applied the One Shot prescription of Bradler and Adami [29] to the first model [1]. The second model begins with the initial state
@@ 3.1 The Basic Black Hole Waterfall Mechanism
with (i) the pump/BH in its vacuum state, (ii) all the interior idlers/Hawking partner particles inside the horizon being in their vacuum date except the $D$-th mode at frequency $\omega_{p}/2^{D}$, and (iii) all the signal modes having occupation number ${n_{p0}}$. Thus, energy distributed all the external signal modes/Hawking radiation is $\sum_{d=1}^{D}({\textstyle\frac{1}{2}})^{d}\,{n_{p0}}\omega_{p}=\left(1-({\textstyle\frac{1}{2}})^{D}\right){n_{p0}}\omega_{p}$, while behind the horizon the pump/BH has zero mass, and we are left with a vanishingly small energy $({\textstyle\frac{1}{2}})^{D}{n_{p0}}\omega_{p}$ in internal idler/Hawking partner particles in mode $i_{D}$. In the limit $D\to\infty$ these energies asymptote to the entire initial mass ${n_{p0}}\omega_{p}$ of the BH appearing the Hawking radiation, and zero mass left in the interior Hawking partner particles. This “waterfall” of cascading idler pumps is qualitatively reminiscent of “ER=EPR” and Replica Wormhole/Island mechanism discussed in the introduction in the sense that both function to move interior Hawking particle particles behind the horizon unitarliy to the external Hawking radiation as the evaporation evolves. However, the BHW mechanism need not invoke wormholes as a physical process. The BHW mechanism’s possible relationship to the Island effect is unclear at this point, though remains an intriguing avenue for further investigation. In the next section we apply the One Shot Trotterizaiton of the second model [2] to this BHW mechanism.
@@ 3.3.1 $\mathbf{D=1:}$
The large spike in the BH temperature $T_{BH}$ (gray dashed curve) around $x_{Page}=163$ occurs just as the the BH entropy $S(\rho_{BH})$ has zero slope, and hence begins to “roll over.” The presence of such a temperature spike at the Page time is already anticipated from the generic Page curve Fig.(1), without assuming any specific BH evaporation model, since from Eq.(61) the denominator $dS_{BH}/dt$ goes to zero at $t_{Page}$ at point where the slope in the numerator $dE_{BH}/dt$ is non-zero (the slope $dE_{BH}/dt\to 0$ only at the end stages where the BH has nearly completely evaporated, $t\to x\geq 200$ in Fig.(8)). After the Page time, there is a rapid (exponential) drop in both the entropy $S_{BH}$ and mass/energy $\bar{E}_{BH}$ of the BH, as revealed in Fig.(8).
@@ 3.3.2 $\mathbf{D=3}$ and $\mathbf{D=4:}$
The other differences between the $D=1$ and $D=(3,4)$ is that the Page Information curve (gray dotdashed curves in Fig.(9)) are (i) essentially zero longer for early times ($0\lesssim x\lesssim 50$), and (ii) rises less steeply for longer times ($50\lesssim x\lesssim 155$) than in the $D=1$ case (gray solid curve inFig.(8)). In all $D$ cases, the Page Information rises exponentially around the crossover $\bar{E}_{BH}=\bar{E}_{R}$ (intersection of the dashed black and dashed red curves), which is expected from the Page Curve. If we were to plot $S_{\textrm{\tiny Bekenstein-Hawking}}\equiv 2\,\pi\,\bar{E}^{2}_{BH}$ (not shown), the cross over in the energies occurs slightly before the crossover in $S_{\textrm{\tiny Bekenstein-Hawking}}=S_{BH}$ [44]. From the Page curve Fig.(1), both crossover conditions could be used to define the Page time $t_{Page}$.
@@ 5 Discussion and Future Work
One advantage of the BH One Shot Waterfall model described in this work is that it does provide a pure state wavefunction (approximation) for the composite BH, (all) signals/idlers $BH\cup S\cup I$ system, from which the von Neumann entropy (and other entropies and relevant quantities) can be computed explicitly. It might be argued that its one major drawback is that it is totally devoid of geometric content (no semiclassical gravitational background is assumed, nor apparent in the its derivation). However, we would argue that the geometry is, in sense, implicit in this model, through the use of the SPDC process. As discussed in the introduction, SPDC and the generation of two-mode squeezed vacuum states (i.e. signal/idler Hawking pairs), appears naturally for accelerated observers in both flat spacetime (Unruh effect), as well as stationary spacetimes (Hawking effect), both with a temporal Killing vector. The squeezing occurs due to the accelerated collapsing surface of the BH horizon, akin to an accelerated mirror [35], and hence is generic in a BH evaporation process, one could argue, essentially independent of the geometry (recall the generation of a squeezed state from a simple harmonic oscillator rapidly changing its frequency under the sudden approximation, discussed on page 10). The goals of the BH One Shot Waterfall model were to capture the essential features of the BH evaporation process, and be able to naturally reproduce a Page curve for the resulting entropies. This model appears to accomplish these goals.
@@
As to any direct connection of the BH One Shot Waterfall model to the modern viewpoint of BH evaporation, one can only speculate at this point. It is curious that the function of the Replica Wormholes and the formation of the interior Island in the modern view is to associate a portion of the interior Hawking partner particles with the external Hawking radiation, so that as time evolves, much of the interior particles are moved to the exterior until the entropy of the BH decreases as the Bekenstein-Hawking formula $S_{\textrm{\tiny Bekenstein-Hawking}}={\textstyle\frac{1}{4}}Area(BH)$. In the BH One Shot Waterfall model a seemingly related functionality is achieved by the idlers/Hawking partner particles acting as subsequent SPDC pump sources. The latter effect is negligible at early times, and only contributes once substantial intermediate population build up in the idlers (all the while they are being simultaneously bled away downstream by their own creation of signal/idler pairs). This appears qualitatively “reminiscent” of the buildup of the Island effect in the modern view, but a direct quantitative comparison is yet to be made (if even possible).

### v2

@@ 2.1 The model of [1]
For physical BHs, a simple argument due to Zurek and Thorne [37] can be used to estimate ${n_{p0}}$ for a BH of initial mass $M_{0}$. Let the BH be composed of ${n_{p0}}$ particles of fixed mass $m_{0}$ of Compton wavelength $\frac{\hbar}{m_{0}c}$ such that $M_{0}={n_{p0}}m_{0}$. These collection of particles of mass $m_{0}$ cannot fit into a size smaller than the Schwarzschild radius $r_{s}=2GM/c^{2}$ of the BH without creating a BH themselves. Thus, setting $\frac{\hbar}{m_{0}c}\sim\frac{2GM_{0}}{c^{2}}$ yields $m_{0}=\frac{1}{2}M_{PL}(\frac{M_{PL}}{M_{0}})$. Setting $M_{0}={n_{p0}}\,m_{0}$ yields ${n_{p0}}=\frac{1}{2}(\frac{M_{PL}}{m_{0}})^{2}\sim\frac{1}{2}(\frac{10^{19}GeV}{m_{0}c^{2}})^{2}$, or $(\frac{m_{0}}{M_{PL}})^{2}=\frac{1}{2\,{n_{p0}}}$. As the BH evaporates, we take its energy to be $E_{BH}=E_{p}(\tau)=\bar{n}_{p}(\tau)m_{0}\,c^{2}=M(\tau)c^{2}$, with $\bar{n}_{p}(0)={n_{p0}}$. The BH thermodynamic (coarse grained) entropy is then given by $S_{\textrm{\tiny Bekenstein-Hawking}}(\tau)=\frac{1}{4}(4\pi r^{2}_{s})=4\pi(\frac{M(\tau)}{M_{PL}})^{2}=4\pi\bar{n}_{p}^{2}(\tau)(\frac{m_{0}}{M_{PL}})^{2}=2\pi\frac{\bar{n}_{p}^{2}(\tau)}{{n_{p0}}}\sim E^{2}_{BH}$  [^1] If one were to allow $m_{0}\to m(\tau)$ to vary with the changing mass of the BH $M(\tau)=\bar{n}_{p}(\tau)\,m(\tau)$, an equivalent calculation yields $(\frac{m(\tau)}{M_{PL}})^{2}=\frac{1}{2\,\bar{n}_{p}(\tau)}$, so that $E_{BH}(\tau)=\bar{n}_{p}(\tau)\,m(\tau)=\frac{1}{\sqrt{2}}\,\sqrt{\bar{n}_{p}(\tau)}\,M_{PL}$, which yields $S_{\textrm{\tiny Bekenstein-Hawking}}(\tau)=2\,\pi\,\bar{n}_{p}(\tau)$, so that once again, $S_{\textrm{\tiny Bekenstein-Hawking}}(\tau)\sim E^{2}_{p}(\tau)$. For simplicity, in this work we will follow the scaling in the main text with fixed mass $m_{0}$, and measure the BH energy as $E_{BH}(\tau)/m_{0}=\bar{n}_{p}(\tau)$., while the fine-grained von Neumann entropy is given by $S(\rho_{BH})\leavevmode\nobreak\ =\leavevmode\nobreak\ S(\rho_{p})\leavevmode\nobreak\ \equiv\leavevmode\nobreak\ S(\rho_{i,s})$ (with the last equality following since the composite state is pure).
@@
Figure 3: (left) Evolution of the average number or particles in the pump/BH $\bar{n}_{p}(\tau)$ (black dashed line), and the signal/Hawking radiation $\bar{n}_{s}(\tau)$ (black solid) for the BH in an initial coherent state with $\bar{n}_{p}(0)$ = 35, and the signal/idlers in the vacuum state. $d^{eff}_{p}\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr{\scriptsize\textrm{def}}\cr\kern 2.0pt\cr=\cr\kern-2.0pt\cr}}}1+\Delta\bar{n}_{p}$ (gray dashed) and $d^{eff}_{s}\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr{\scriptsize\textrm{def}}\cr\kern 2.0pt\cr=\cr\kern-2.0pt\cr}}}1+\Delta\bar{n}_{s}$ (gray solid) are the effective dimensions of the pump/BH and signal/Hawking radiation in terms of the variance in their particle number [1, 30]. Vertical red-dashed line indicates the time at which $d\bar{n}_{p}(\tau)/d\tau=0$ again for $\tau>0$, the limit of the validity of the model. (right) von Neumann entropies for: $S(\rho_{s})$ signal/Hawking radiation (dashed-black curve), $S(\rho_{thermal})$ effective thermal state (solid black curve) formed from $\rho_{s}$. The difference between the solid-black and dashed-black curve is the Page Infomation $I(\tau)\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr{\scriptsize\textrm{def}}\cr\kern 2.0pt\cr=\cr\kern-2.0pt\cr}}}S(\rho_{thermal})-S(\rho_{s})$.
In Fig.(3)(left) we plot the evolution of the average number or particles $\bar{n}_{p}(\tau)$ in the pump/BH (black dashed curve), and in the signal/Hawking radiation $\bar{n}_{s}(\tau)$ (black solid), for the BH in an initial coherent state with $\bar{n}_{p}(0)=35=\alpha^{2}$  [^2] The plots of $\bar{n}_{p}(\tau)$ and $\bar{n}_{s}(\tau)$ have a similar form for any value of ${n_{p0}}$ as long as it significantly different than the initial number of particles in the signal/idler modes, which we have taken to be zero. Plots using large initial values of ${n_{p0}}$ simply change the scale on the $\tau$ axis., and the signal/idlers in the vacuum state, $|\psi(0)\rangle=|\alpha\rangle_{p}|0\rangle_{i}|0\rangle_{s}$, with $|\alpha\rangle_{p}=e^{-\alpha^{2}/2}\sum_{m=0}^{\infty}\frac{\alpha^{m}}{\sqrt{m!}}|m\rangle_{p}$ [4, 5]. For early times both $\bar{n}_{p}(\tau)$ and $\bar{n}_{s}(\tau)$ have nearly zero slope, and for these time with $\bar{n}_{p}(\tau)\gg\bar{n}_{s}(\tau)$ the Hawking radiation is essentially composed of squeezed signal/idler pairs, and hence is thermal. We define an effective thermal state from $\rho_{s}$ by defining $\rho_{thermal}=\sum_{m=0}^{\infty}\frac{\bar{n}_{s}^{m}(\tau)}{(\bar{n}_{s}(\tau)+1)^{m+1}}|m\rangle_{s}\langle m|$. The Page Information [30, 1] $I(\tau)\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr{\scriptsize\textrm{def}}\cr\kern 2.0pt\cr=\cr\kern-2.0pt\cr}}}S(\rho_{thermal})-S(\rho_{s})$ represents the information that leaks out of the BH due to the correlations (entanglement) built up between the pump/BH and signal/Hawking radiation as the system evolves. This is shown in Fig.(3)(right).
@@
We will see in the third model considered in Section 3, constructed in order to address model objective (3), this issue can also be rectified by assuming that the interior Hawking particles can also create their own interior/exterior idler/signal Hawking pairs, now with half of each interior particle’s energy, i.e. $\omega_{p}/4$. With each newly created idler acting as its own “pump”, this cascading (waterfall) process can be carried out ad infinitum so that the total energy carried away by the collection of all the signals/Hawking radiation is ${n_{p0}}\,\omega_{p}\,\sum_{d=1}^{D}(\frac{1}{2})^{d}=(1-(\frac{1}{2})^{D})\,{n_{p0}}\,\omega_{p}\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr{\scriptstyle{{D\to\infty}}}\cr\kern 2.0pt\cr{=\;\;}\cr\kern-2.0pt\cr}}}{n_{p0}}\,\omega_{p}$, the initial mass of the BH (this “remnant mitigation” proposal was first suggested by the author in [41]). What remains inside the horizon are ${n_{p0}}$ idler particles of arbitrarily small vanishing energy $(\frac{1}{2})^{D}\,{n_{p0}}\,\omega_{p}\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr{\scriptstyle{{D\to\infty}}}\cr\kern 2.0pt\cr{=\;\;}\cr\kern-2.0pt\cr}}}0$. This phenomena of zero energy remaining in the internal (“soft-hair”) horizon modes of the BH was explored in a 2018 paper by Hotta, Nambu and Yamaguchi entitled Soft-hair-enhanced entanglement beyond the Page Curves in Black Hole evaporation Qubit model [3]. In this work, an infinite number of super-translation charges appear to be able to store the whole quantum information of the absorbed matter forming the BH, arising from a near-horizon symmetry providing a massive degeneracy of the Hamiltonian (the approximately zero energy “soft hair” of the BH horizon). The impetuous for the BH Waterfall Model presented here in this work was inspired by this work, with a desire to develop a qualitatively similar model, but from a squeezing Hamiltonian perspective. Before we explore this model in Section 3, which we descriptively denote as the Black Hole Waterfall, we first review the second One Shot model [2] in the next section, on which the former is built.
@@ 2.2 The One Shot model of [2]
The second model [2] applied the One Shot prescription of Brádler and Adami [29] to the first model [1]. The second model begins with the initial state
@@ 3.1 The Basic Black Hole Waterfall Mechanism
with (i) the pump/BH in its vacuum state, (ii) all the interior idlers/Hawking partner particles inside the horizon being in their vacuum date except the $D$-th mode at frequency $\omega_{p}/2^{D}$, and (iii) all the signal modes having occupation number ${n_{p0}}$. Thus, energy distributed all the external signal modes/Hawking radiation is $\sum_{d=1}^{D}({\textstyle\frac{1}{2}})^{d}\,{n_{p0}}\omega_{p}=\left(1-({\textstyle\frac{1}{2}})^{D}\right){n_{p0}}\omega_{p}$, while behind the horizon the pump/BH has zero mass, and we are left with a vanishingly small energy $({\textstyle\frac{1}{2}})^{D}{n_{p0}}\omega_{p}$ in internal idler/Hawking partner particles in mode $i_{D}$. In the limit $D\to\infty$ these energies asymptote to the entire initial mass ${n_{p0}}\omega_{p}$ of the BH appearing the Hawking radiation, and zero mass left in the interior Hawking partner particles. This “waterfall” of cascading idler pumps is qualitatively reminiscent of “ER=EPR” and Replica Wormhole/Island mechanism discussed in the introduction in the sense that both function to move interior Hawking particle particles behind the horizon unitarily to the external Hawking radiation as the evaporation evolves. However, the BHW mechanism need not invoke wormholes as a physical process. The BHW mechanism’s possible relationship to the Island effect is unclear at this point, though remains an intriguing avenue for further investigation. In the next section we apply the One Shot Trotterization of the second model [2] to this BHW mechanism.
@@ 3.3.1 $\mathbf{D=1:}$
The large spike in the BH temperature $T_{BH}$ (gray dashed curve) around $x_{Page}=163$ occurs just as the BH entropy $S(\rho_{BH})$ has zero slope, and hence begins to “roll over.” The presence of such a temperature spike at the Page time is already anticipated from the generic Page curve Fig.(1), without assuming any specific BH evaporation model, since from Eq.(61) the denominator $dS_{BH}/dt$ goes to zero at $t_{Page}$ at point where the slope in the numerator $dE_{BH}/dt$ is non-zero (the slope $dE_{BH}/dt\to 0$ only at the end stages where the BH has nearly completely evaporated, $t\to x\geq 200$ in Fig.(8)). After the Page time, there is a rapid (exponential) drop in both the entropy $S_{BH}$ and mass/energy $\bar{E}_{BH}$ of the BH, as revealed in Fig.(8).
@@ 3.3.2 $\mathbf{D=3}$ and $\mathbf{D=4:}$
The other differences between the $D=1$ and $D=(3,4)$ is that the Page Information curve (gray dotdashed curves in Fig.(9)) are (i) essentially zero longer for early times ($0\lesssim x\lesssim 50$), and (ii) rises less steeply for longer times ($50\lesssim x\lesssim 155$) than in the $D=1$ case (gray solid curve inFig.(8)). In all $D$ cases, the Page Information rises exponentially around the crossover $\bar{E}_{BH}=\bar{E}_{R}$ (intersection of the dashed black and dashed red curves), which is expected from the Page Curve. If we were to plot $S_{\textrm{\tiny Bekenstein-Hawking}}\equiv 2\,\pi\,\bar{E}^{2}_{BH}$ (not shown), the cross over in the energies occurs slightly before the crossover in $S_{\textrm{\tiny Bekenstein-Hawking}}=S_{BH}$  [^5] For $D=3$, the energy crossover $\bar{E}_{BH}=\bar{E}_{R}$ and entropy crossover $S_{\textrm{\tiny Bekenstein-Hawking}}=S_{BH}$ occur at the approximately the same temporal point $x$ if $S_{\textrm{\tiny Bekenstein-Hawking}}$ were defined instead as $S_{\textrm{\tiny Bekenstein-Hawking}}\approx(2\,\pi/3)\,\bar{E}^{2}_{BH}$. From the Page curve Fig.(1), both crossover conditions could be used to define the Page time $t_{Page}$.
@@ 5 Discussion and Future Work
One advantage of the BH One Shot Waterfall model described in this work is that it does provide a pure state wavefunction (approximation) for the composite BH, (all) signals/idlers $BH\cup S\cup I$ system, from which the von Neumann entropy (and other entropies and relevant quantities) can be computed explicitly. It might be argued that its one major drawback is that it is totally devoid of geometric content (no semiclassical gravitational background is assumed, nor apparent in the its derivation). However, we would argue that the geometry is, in sense, implicit in this model, through the use of the SPDC process. As discussed in the introduction, SPDC and the generation of two-mode squeezed vacuum states (i.e. signal/idler Hawking pairs), appears naturally for accelerated observers in both flat spacetime (Unruh effect), as well as stationary spacetimes (Hawking effect), both with a temporal Killing vector. The squeezing occurs due to the accelerated collapsing surface of the BH horizon, akin to an accelerated mirror [35], and hence is generic in a BH evaporation process, one could argue, essentially independent of the geometry (recall the generation of a squeezed state from a simple harmonic oscillator rapidly changing its frequency under the sudden approximation, discussed on page 10). Another advantage of models based on a trilinear Hamiltonian, is that they hold the prospect of being experimentally testable, as evidenced by the recent thesis of Gutiérrez [42], which was directly inspired by the work of Brádler and Adami [29]. The goals of the BH One Shot Waterfall model were to capture the essential features of the BH evaporation process, and be able to naturally reproduce a Page curve for the resulting entropies. This model appears to accomplish these goals.
@@
As to any direct connection of the BH One Shot Waterfall model to the modern viewpoint of BH evaporation, one can only speculate at this point. One of the central questions of black hole evaporation is the question of where exactly do the degrees of freedom reside that make up $S_{\textrm{\tiny Bekenstein-Hawking}}$. One possible view (see the review by D.N. Page [43]) is that in a full theory of quantum gravity one might not expect a definite four-metric, or even a definite causal structure. This could in turn blur the absolute distinction between the inside and outside of the BH. There still remains the difficult question of how the information of those degrees of freedom can get out as the BH evaporates. It is curious that the function of the Replica Wormholes and the formation of the interior Island in the modern view is to associate a portion of the interior Hawking partner particles with the external Hawking radiation, so that as time evolves, much of the interior particles are moved to the exterior until the entropy of the BH decreases as the Bekenstein-Hawking formula $S_{\textrm{\tiny Bekenstein-Hawking}}={\textstyle\frac{1}{4}}Area(BH)$. In the BH One Shot Waterfall model a seemingly related functionality is achieved by the idlers/Hawking partner particles acting as subsequent SPDC pump sources. The latter effect is negligible at early times, and only contributes once substantial intermediate population build up in the idlers (all the while they are being simultaneously bled away downstream by their own creation of signal/idler pairs). This appears qualitatively “reminiscent” of the buildup of the Island effect in the modern view, but a direct quantitative comparison is yet to be made (if even possible).
While a full theory of quantum gravity remains as of yet elusive, the robust features of the semiclassical BH evaporation calculation by Hawking (and Unruh) should emerge in the low energy limit. Since the process of squeezing is at the heart of these semiclassical calculations, the bilinear Hamiltonian of the optical SPDC process serves as reasonable phenomenological approximation when the BH mass is either fixed, or considered so large as essentially to be constant [28]. The next level of approximation to attempt to incorporate some level of back-reaction in the BH evaporation process, would be to allow the mass of the BH to decrease. The trilinear model involved here, and studied by previous authors [1, 2, 29, 30, 42] is the simplest, and most natural Hamiltonian extension of the prior bilinear SPDC model. One could reasonably argue that a low-energy approximation of any effective Hamiltonian arising from a consistent theory of quantum gravity must incorporate some form of a tripartite interaction between the BH, the emitted Hawking radiation and their infalling partner particles, and hence would be most simply phenomenologically modeled by a trilinear Hamiltonian of SPDC with a depleting “pump.” Such low energy calculations would be kind of calculations that would first be performed if a consistent theory of quantum gravity existed.